The difference between the measures of two complementary angles is 58 degrees. What is the measure of the two angles if using two variables (x and y).
Tricia

Answer by prabhjyot(165)

complementary angles are the angles which have sun equal to 90
Let two angles be x and y
so x+y=90 ----->(1)
given that difference of the two numbers is 58
that means x-y=58 ----->(2)
from equation (2) x=58+y
now putting the value of x in equation in (1)
58+y+y=90
58+2y=90
2y=90-58
y=16
now plug the value of x in either of the two equation
we get x=74
Therefore two angles are x=74 and y=16